The sum of three numbers is 134.The second number is 9 less than 3 times the first number.The third number is 5 more than 2 times the first number.Find the three numbers
Answer by elima(1433)

Let x be the first number
3x-9; second number
2x+5; third number
x + 3x-9+2x+5=134
collect like terms;
x+3x+2x-9+5=134
6x-4=134
6x=138
x=23
so; our first number is 23
3(23)-9=60; second number
2(23)+5=51; third number
23+60+51=134
:)